# Tunable lower critical fractal dimension for a non-equilibrium phase transition

### v1

@@ VI Time evolution of the density profile
Additional insight can be gained by examining the time evolution of the density profile. In particular, we focus on the spatial structure of the slowest decaying mode in the fractal geometry corresponding to the Sierpinski carpet ($d_{H}\approx 1.8$). Figure S.5 displays the spatial dependence of the density deviation from the steady state at different time steps; specifically, we plot $|n_{ss}(\textbf{r})-n(\textbf{r},t)|$ for various times $t$. This local information can be compared with the global behavior shown in the upper panel of Fig. S.5, where the evolution of the spatially averaged density $\bar{n}(t)$ is presented for the same system.
@@
Figure S.5: Lower panels: Spatial profile of the density deviation from the steady state at different time steps $t$, given by $|n_{ss}(\textbf{r})-n(\textbf{r},t)|$, for the Sierpinski carpet at the crossover point $I=2.3\gamma^{2}$ with $\Delta=\gamma$. Upper panel: Time evolution of the spatially averaged density $\bar{n}(t)=\int\mathrm{d}^{2}x\,n(t,x)$. In red, we plot the logarithm of the deviation from the steady state, making the asymptotic decay rate clearly visible.
@@
At early times, the density $n(\textbf{r},t)$ is close to zero throughout the system, so the difference shown in the first lower panel of Fig. S.5 primarily reflects the shape of the final steady state. At later times, as shown in the final panels, a distinct pattern begins to emerge. This structure corresponds to the longest-lived mode—i.e., the one governing the asymptotic decay rate. In the Liouvillian framework, this is the eigenstate of the Lindbladian superoperator associated with the Liouvillian gap (the smallest nonzero eigenvalue). Thus, we are observing the mode that is key to characterizing the system’s critical behavior.
@@
For a homogeneous condensate, this connects to the dispersion relation $\omega(\textbf{k})$ given in Eq. (S.2). In certain cases, the slowest-decaying mode (i.e., the one with the largest imaginary part of $\omega$) occurs at a nonzero wavevector $\textbf{k}\neq\vec{0}$, meaning that the longest-lived mode may not share the spatial symmetry of the driving field. In our case, the resulting mode exhibits a distinctive and nontrivial spatial structure, which will be the subject of future investigation.
@@
VII Critical Detuning Independence from the interaction strength within the Non-Equilibrium Gross-Pitaevskii Framework
@@
$i\frac{d\psi_{2}}{dt}=\left[-\Delta-\frac{\hbar\nabla^{2}}{2m}+\frac{g_{1}}{\eta^{2}}|\psi_{2}|^{2}-i\frac{\gamma}{2}\right]\psi_{2}+iF_{2}(\textbf{r}).$ (S.5)

### v2

@@ VII Time evolution of the density profile
Additional insight can be gained by examining the time evolution of the density profile. In particular, we focus on the spatial structure of the slowest decaying mode in the fractal geometry corresponding to the Sierpinski carpet ($d_{H}\approx 1.8$). Figure S8 displays the spatial dependence of the density deviation from the steady state at different time steps; specifically, we plot $|n_{ss}(\textbf{r})-n(\textbf{r},t)|$ for various times $t$. This local information can be compared with the global behavior shown in the upper panel of Fig. S8, where the evolution of the spatially averaged density $\bar{n}(t)$ is presented for the same system.
@@
Figure S8: Lower panels: Spatial profile of the density deviation from the steady state at different time steps $t$, given by $|n_{ss}(\textbf{r})-n(\textbf{r},t)|$, for the Sierpinski carpet at the crossover point $I=2.3\gamma^{2}$ with $\Delta=\gamma$. Upper panel: Time evolution of the spatially averaged density $\bar{n}(t)=\int\mathrm{d}^{2}x\,n(t,x)$. In red, we plot the logarithm of the deviation from the steady state, making the asymptotic decay rate clearly visible.
@@
At early times, the density $n(\textbf{r},t)$ is close to zero throughout the system, so the difference shown in the first lower panel of Fig. S8 primarily reflects the shape of the final steady state. At later times, as shown in the final panels, a distinct pattern begins to emerge. This structure corresponds to the longest-lived mode—i.e., the one governing the asymptotic decay rate. In the Liouvillian framework, this is the eigenstate of the Lindbladian superoperator associated with the Liouvillian gap (the smallest nonzero eigenvalue). Thus, we are observing the mode that is key to characterizing the system’s critical behavior.
@@
For a homogeneous condensate, this connects to the dispersion relation $\omega(\textbf{k})$ given in Eq. (S2). In certain cases, the slowest-decaying mode (i.e., the one with the largest imaginary part of $\omega$) occurs at a nonzero wavevector $\textbf{k}\neq\vec{0}$, meaning that the longest-lived mode may not share the spatial symmetry of the driving field. In our case, the resulting mode exhibits a distinctive and nontrivial spatial structure, which will be the subject of future investigation.
@@
VIII Critical Detuning Independence from the interaction strength within the Non-Equilibrium Gross-Pitaevskii Framework
@@
$i\frac{d\psi_{2}}{dt}=\left[-\Delta-\frac{\hbar\nabla^{2}}{2m}+\frac{g_{1}}{\eta^{2}}|\psi_{2}|^{2}-i\frac{\gamma}{2}\right]\psi_{2}+iF_{2}(\textbf{r})~.$ (S5)